A density bottle contains 104.5 degree centigrade of a liquid at 34 degree centigrade on heating it to 100 degree centigrade the mass of the liquid expelled is 4.5gms.Then find the coefficient of apparent expansion of liquid.

Grade:11

3 Answers

Nirmal Singh.
10 years ago
Let the density of the liquid in the density bottle be `d`
Then volume of the liquid at 34degree centigrade V1=md=104.5d
volume of the liquid expelled when it is heated to 100degree centigrade(V2-V1)=4.5d
Coefficient of apparent expansion of liquid=[V2-V1]/[V1*(100-34)]
=[4.5d]/[104.5d*66]
=0.000652458
Regards,
Nirmal Singh
Askiitians faculty
=0.000652458
=0.000652458
sivakrishna
10 years ago
let the density of the liquid in the bottle be d. Given , m1=104.5gms t1=34 degree centigrade m2=m1+4.5 =104.5+4.5 =109gms t2=100degree centigrade we know that the coefficient of apparent expansion=(V2-V1)/V1(t2-t1) d=m/V>v=m/d = (m2/d-m1/d)m1/d*(100-34) =(109/d-104.5/d)/104.5/d*(66) =(4.5/d)/6897/d =4.5/6897 =0.000652457 =0.0007/degree centigrade
Vyshnavi
10 years ago
Let density of liquid in the liquid bottle be `d` Then volume of liquid at 34 degree centigrade v1=md=104.5d Volume of liquid expelled when it is heated to 100 degree centigrade(V2-V1)=4.5d Coefficient of apparent expansion of liquid =[V2-V1]/[V1*(100-34)] =[4.5d]/[104.5d*66] =0.000652457 =0.0007/degree centigrade